OpenStudy (anonymous):

(5/7)3 3 is a exponent solve it

OpenStudy (anonymous):

\[\left( \frac{ 5 }{ 7 } \right)^{3} = \frac{ 5^{3} }{ 7^{3} }\] good luck solving it

OpenStudy (anonymous):

Thank you !!!!!!!